# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ Abstract.
The space of chains on a compact connected space encodes all the different ways of continuously growing out of a point until exhausting the space. A chain is generic if its orbit under the action of the underlying homeomorphism group is comeager: a space $X$ has a generic chain if there is essentially just one type of chain on $X$. Gutman, Tsankov and Zucker proved that compact manifolds of dimension at least 3 do not have a generic chain. We extend and generalize their result, covering a large class of spaces which includes all compact surfaces except for the sphere and the real projective plane – for which the question remains open – as well as all other homogeneous Peano continua, circle excluded. If the spaces are moreover strongly locally homogeneous, which is the case for any closed manifold as well as the Menger curve, we prove that chains cannot be classified up to homeomorphism by countable structures, and that the underlying homomorphism groups have non-metrizable universal minimal flows. This is in contrast with the case of $1$-dimensional manifolds: their homomorphism groups have metrizable universal minimal flow and we show that their chains are classifiable by countable structures and have a generic element. The proof of the main result relies on crafting a dictionary between open sets of chains on one side, and walks on finite connected graphs on the other. We then establish a novel combinatorial necessary condition for the existence of a generic chain, an off-by-one weak amalgamation principle, and prove that it is not satisfied under the hypotheses of our theorem.
@@
G.B. was partially supported by Istituto Nazionale di Alta Matematica "Francesco Severi". A.C. and A.V. were supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044–-390685587, Mathematics Münster: Dynamics–Geometry–Structure, through SFB 1442. A.C. was moreover supported by the ERC Starting Grant 101077154 “Definable Algebraic Topology”. A.V. was moreover supported by the ERC Advanced Grant 834267–AMAREC
@@ 1. Introduction
Given a metrizable compact connected space $X$, a maximal chain of compact connected sets – just chain, from now on –, is a collection of compact connected subsets of $X$ which is linearly ordered by inclusion and maximal with respect to this property. Equivalently, it is a homeomorphic image of the interval $[0,1]$ in the hyperspace $C(X)$ of compact connected subsets of $X$ such that 00 is mapped to a singleton, $1$ to the point $X\in C(X)$ and for each $s<t$, the image of $s$ is a subset of the image of $t$ [illanes1999hyperspaces]*Lemma 14.7. The collection $\Phi(X)$ of all such chains therefore represents all the different ways one can start, say at time 00, from a point of $X$, and then continuously grow out of it, until, at time $1$, they have exhausted $X$.
We are therefore interested in classifying such chains up to homeomorphisms of $X$, but this is only feasible for very simple spaces, such as the the interval or the circle, see Appendix A. At the very least, we can ask for which spaces there is essentially only one chain, that is, whether there is $\mathcal{C}\in\Phi(X)$ such that its orbit under the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$ is comeager in $\Phi(X)$, endowed with its natural compact hyperspace topology. In such case, we say that $\mathcal{C}$ is a generic chain.
We show that the chain on the circle consisting of all intervals centered around some point $x_{0}$ is generic in $\Phi(S^{1})$, but Gutman, Tsankov and Zucker proved in [GTZ] that no closed manifold of dimension at least $3$ has generic chains, and nor does the Hilbert cube. The case of surfaces was left open; a corollary of our main theorem is that
@@
Our theorem concerns a vastly larger class of spaces, of all positive dimensions, and recovers the aforementioned result of [GTZ]. A Peano continuum is a metrizable compact connected space which is locally connected. Among notable examples which are not manifolds are the Sierpiński carpet and Menger curve. These are universal one-dimensional spaces obtained by carving out the square and the cube, respectively.
A subset $K\in X$ is locally separating if $U\setminus K$ is disconnected for some connected open $U\subseteq X$. A space is planar if it admits an embedding in $\mathbb{R}^{2}$ and it is locally non-planar if none of its nonempty open sets are planar. Let us say a Peano continuum $X$ has a circular covering if it admits a covering $O_{0},\dots,O_{\ell-1}$ in $\ell\geq 4$ connected open subsets such that $\operatorname{Cl}(O_{i})\cap\operatorname{Cl}(O_{j})\neq\emptyset$ if and only if $\lvert i-j\rvert\leq 1\text{ mod}\,\ell$, and $O_{i}\setminus\bigcup_{j\neq i}O_{j}$ is connected for each $i<\ell$.
@@
Let us parse the above hypotheses for $X$ a compact manifold. A compact manifold has a locally separating point if and only if it is $1$-dimensional, and it has a locally non-planar open subset if and only if it is at least $3$-dimensional. A boundary of a surface is a simple closed curve which is not locally separating and lives in a planar open set. Finally, the sphere and the real projective plane are the only closed surfaces without a circular covering, see Proposition 6.13.
@@
We prove Theorem 1.2 by first establishing a dictionary between Peano continua and open sets of chains on one side, and finite connected graphs and walks on these graphs, on the other. We find a combinatorial necessary condition for the existence of a generic chain on a Peano continuum $X$, and prove that it is does not hold under our assumptions on $X$.
@@
The combinatorial condition is an off-by-one weak amalgamation principle, roughly stating that each walk can be refined by another walk in such a way that any two different walks refining it can be reconciled by a small perturbation. We extract the combinatorial idea at the core of the proof in [GTZ] – that chains can be made irreconcilable by winding enough times in opposite directions – and discretize it, producing irreconcilable walks which wind around circular subgraphs.
@@
The discretization relies on a handful of classical continuum-theoretical black-boxes, all of which are stated in Section 4. In the following two sections, these black-boxes are used to derive basic combinatorial statements about graphs and walks on graphs. The proof of Theorem 1.2 is purely combinatorial.
@@
This approach to the study of topological and dynamical properties of compact metrizable spaces by combinatorial means traces its roots to the field of projective Fraïssé theory, first introduced in [Irwin2006]. In rough terms, one constructs an infinite compact zero-dimensional graph, the combinatorial model of $X$, and finite approximations of such graph, and studies them to derive properties about $X$, leaning on the dictionary between the combinatorics of countable structures and the dynamics of Polish non-archimedean groups put forth in the seminal paper [KPT]. The aforementioned Menger curve was studied under this lens in [PS22]. It follows from the results in [BK:Lelek], obtained in a projective Fraïssé theoretical framework, that the Lelek fan, a one-dimensional continuum which is not locally connected, has a generic downward closed chain.
@@
Nonetheless, our strategy is innovative in that it is the first instance in which non-trivial necessary conditions for the existence of a comeager orbit have been translated to combinatorial statements. In this paper, we adopt a minimalistic approach and we restrict ourselves to only finite combinatorial objects, though our proofs could also be used to derive the analogous statements for the combinatorial models of the spaces involved.
@@
Theorem 1.2 has important dynamical consequences for the groups of homomorphisms involved. If $G$ is a topological group, we call a continuous action $G\curvearrowright Y$ on a compact space a $G$-flow. A $G$-flow is minimal if every orbit is dense. Both classifying minimal flows of a group $G$ and, viceversa, classifying groups based on the properties of their minimal flows are central goals in topological dynamics. Some groups of homeomorphisms have very simple minimal dynamics: the only minimal flow of the group $\operatorname{Homeo}_{+}([0,1])$ of orientation preserving homeomorphisms of the interval is the trivial action and the only non-trivial minimal flow of $\operatorname{Homeo}_{+}(S^{1})$ is the action on the circle itself (see [Pestov]).
@@
A number of dividing lines regarding the complexity of the minimal dynamics of a topological group have been studied, and are particularly well understood when $G$ is Polish, i.e., separable and completely metrizable. A Polish group is extremely amenable if all of its minimal flows are trivial, it is CAP if they are metrizable (see [CAP]), and it has the generic point property if every minimal flow has a comeager orbit (see [AKL].) Every extremely amenable group is CAP, every Polish CAP group has the generic point property (see [ben2017metrizable]), and no locally compact group has the generic point property. Such properties can also be stated in terms of the universal minimal flow of $G$: CAP groups are also known as groups with metrizable universal minimal flow.
@@
Until recently (see [kaleido]), there was essentially only one example of Polish group with the generic point property which is not CAP (see [kwiat]). In [GTZ] it is proved that the subgroup $\operatorname{Homeo}_{0}(X)$ of homeomorphism isotopic to the identity is not CAP, for any closed surface $X$, but it is left open whether such groups have the generic point property. Since for closed surfaces $\operatorname{Homeo}_{0}(X)\curvearrowright\Phi(X)$ is minimal, Theorem 1.2 allows us to answer in the negative in all but two cases.
@@
Let us end this introduction with a series of corollaries of the Theorem 1.2. A space $X$ is homogeneous if the action $\operatorname{Homeo}(X)\curvearrowright X$ is transitive. Homogeneous Peano continua are either the circle, or a closed surface, or are locally non-planar (see [MR2674033]*Remark 8.)
@@
On the other hand, the Sierpiński curve, which is the unique one-dimensional planar Peano continuum without locally separating points (see [whyburn]), is not homogeneous. Nonetheless, it has plenty of non locally separating simple closed curves, as it can be constructed as the complement, in the sphere, of countably many pairwise disjoint open disks whose union is dense. In particular, any open set is planar and contains a non locally separating simple closed curve.
A one-dimensional Peano continuum with no locally separating points is either the Menger curve or has a nonempty open connected subset homeomorphic to an open subset of the Sierpiński curve (see [MR1314981]*Lemma 1.1.) Therefore
@@
We prove in Theorem 7.7 that a sufficient condition for minimality of the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$, when $X$ is a Peano continuum without locally separating points, is that $X$ be strongly locally homogeneous, see Section 7 for the definition. For example, all closed manifolds are strongly locally homogeneous. Since these spaces are in particular homogeneous, they have no generic chain, by Corollary 1.5, except possibly if they are the sphere or the real projective plane. Therefore:
@@
If $X$ is a strongly locally homogeneous Peano continuum which is not the circle, the sphere, or the real projective plane, then $\operatorname{Homeo}(X)$ does not have the generic point property. In particular, its universal minimal flow is not metrizable.
Among other examples of spaces which fit the hypotheses of Corollary 1.7 are the Menger curve, the universal $k$-dimensional compactum $\mu^{k}$, and any compact connected $\mu^{k}$-manifold, for $k\geq 1$ (see [MR0920964]*Theorem 3.2.2.)
For the spaces in Corollary 1.7, we are also able to prove that the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$ is turbulent, which implies strong non-classifiability for the orbit equivalence relation of the action. We say that chains on $X$ are classifiable by countable structures if there is a Borel function $\mathcal{M}\colon\Phi(X)\to\mathrm{Mod}(\mathcal{L})$ to the space of structures with domain $\mathbb{N}$ in some countable relational language $\mathcal{L}$ such that
$\mathcal{C}^{\prime}\in\operatorname{Homeo}(X)\cdot\mathcal{C}\Longleftrightarrow\mathcal{M}(\mathcal{C})\approx\mathcal{M}(\mathcal{C}^{\prime}),$
@@
We remark that the first question in 1.4 could be answered in the positive by proving that chains on the sphere are classifiable by countable structures, but we believe that this is not the case.
In Appendix B, we prove that $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ is minimal, for the Sierpiński carpet $S$, and thus that $\operatorname{Homeo}(S)$ does not have the generic point property, and its universal minimal flow is not metrizable. The proof is rather lengthy and technical; it could foreseeably be simplified by finer results on the degree of transitivity of the diagonal actions $\operatorname{Homeo}(S)\curvearrowright S^{n}$, which we were not able to find in the literature. We leave the question of classifiability by countable structures of chains on the Sierpiński carpet open.
@@ 1.1. Plan of the paper
In section 5 we construct the dictionary between open sets of chains of Peano continua and walks on graphs, and prove the required discretization results to establish the combinatorial necessary condition for the existence of a generic chain, which is Theorem 5.8. Section 6 is the heart of the paper and contains the proof of Theorem 1.2, which we obtain as a sum of Theorems 6.4, 6.6, 6.10, and 6.12, whose proofs are completely combinatorial. Section 7 is devoted to the question of minimality and turbulence of $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$, and their consequences for classifiability of chains and topological dynamics of $\operatorname{Homeo}(X)$.

### v3

@@ Abstract.
The space of chains on a compact connected space encodes all the different ways of continuously growing out of a point until exhausting the space. A chain is generic if its orbit under the action of the underlying homeomorphism group is comeager. In this paper we show that a large family of topological spaces do not have a generic chain: in addition to all manifolds of dimension at least 3, for which the result was already known, our theorem covers all compact surfaces except for the sphere and the real projective plane – for which the question remains open – as well as all other homogeneous Peano continua, circle excluded. If the spaces are moreover strongly locally homogeneous, which is the case for any closed manifold and the Menger curve, we prove that chains cannot be classified up to homeomorphism by countable structures, and that the underlying homeomorphism groups have non-metrizable universal minimal flows, with all orbits meager, in contrast to the case of 1-dimensional manifolds. The proof of the main result is of combinatorial nature, and it relies on the creation of a dictionary between open sets of chains on one side, and walks on finite connected graphs on the other.
@@
G.B. was partially supported by Istituto Nazionale di Alta Matematica “Francesco Severi”. A.C. and A.V. were supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044–-390685587, Mathematics Münster: Dynamics–Geometry–Structure, through SFB 1442. A.C. was moreover supported by the ERC Starting Grant 101077154 “Definable Algebraic Topology”. A.V. was moreover supported by the ERC Advanced Grant 834267–AMAREC
@@ 1. Introduction
Given a metrizable compact connected space $X$, a maximal chain of compact connected sets of $X$ – just chain, from now on –, is a collection of compact connected subsets of $X$ which is linearly ordered by inclusion and maximal with respect to this property. Equivalently, it is a homeomorphic image of the interval $[0,1]$ in the hyperspace $C(X)$ of compact connected subsets of $X$ such that 00 is mapped to a singleton, $1$ to the point $X\in C(X)$ and for each $s<t$, the image of $s$ is a subset of the image of $t$ [illanes1999hyperspaces]*Lemma 14.7. The collection $\Phi(X)$ of all such chains therefore represents all the different ways one can start, say at time 00, from a point of $X$, and then continuously grow out of it, until, at time $1$, they have exhausted $X$.
A classification of such chains up to homeomorphisms of $X$ is only feasible for very simple spaces, such as the interval or the circle (see Appendix A). At the very least, we can ask for which spaces there is essentially only one chain, that is, whether there is $\mathcal{C}\in\Phi(X)$ such that its orbit under the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$ is comeager in $\Phi(X)$, endowed with its natural compact hyperspace topology. In such case, we say that $\mathcal{C}$ is a generic chain.
Focusing on compact manifolds, the chain on the circle $S^{1}$ consisting of all intervals centered around some point $x_{0}$ is generic in $\Phi(S^{1})$, and a similar statement holds for the closed interval (see Proposition A.4 and Theorem A.6). On the other hand, Gutman, Tsankov and Zucker proved in [GTZ] that no closed manifold of dimension at least $3$ has generic chains, and nor does the Hilbert cube. The case of surfaces was left open. A corollary of our main theorem gives the following.
@@
We obtain Theorem 1.1 as consequence of a stronger statement which not only recovers the aforementioned result of [GTZ], but which moreover covers a vast class of Peano continua going well beyond the context of manifolds.
A Peano continuum is a metrizable compact connected space which is locally connected. Among notable examples which are not manifolds are the Sierpiński carpet and the Menger curve. These are universal one-dimensional continua obtained by iteratively carving out the square and the cube, respectively, similarly to how the Cantor space is obtained from the unit interval.
Before stating our main theorem, we introduce some basic terminology. A subset $K\subseteq X$ is locally separating if $U\setminus K$ is disconnected for some connected open $U\subseteq X$. A space is planar if it admits an embedding in $\mathbb{R}^{2}$, and it is locally non-planar if none of its nonempty open sets are planar. Recall that a simple closed curve in $X$ is the image of an embedding of the circle $S^{1}$ into $X$. Let us say that a Peano continuum $X$ has a circular covering if it admits a covering $O_{0},\dots,O_{\ell-1}$ in $\ell\geq 4$ connected open subsets such that $\operatorname{Cl}(O_{i})\cap\operatorname{Cl}(O_{j})\neq\emptyset$ if and only if $\lvert i-j\rvert\leq 1\text{ mod}\,\ell$, and $O_{i}\setminus\bigcup_{j\neq i}O_{j}$ is connected for each $i<\ell$.
@@
Let us parse the above hypotheses for $X$ a compact manifold (we refer to [MR2954043] for background and definitions). A compact manifold has a locally separating point if and only if it is $1$-dimensional, and it has a locally non-planar open subset if and only if it is at least $3$-dimensional. A boundary of a surface is a simple closed curve which is not locally separating and lies in a planar open set. Finally, the sphere and the real projective plane are the only closed surfaces without a circular covering (see Proposition 6.13).
@@
We prove Theorem 1.2 by first establishing a correspondence between Peano continua and open sets of chains on one side, and finite connected graphs and walks on these graphs on the other. We then find a combinatorial necessary condition for the existence of a generic chain on a Peano continuum (Theorem 5.9), and prove that it is does not hold under our assumptions.
@@
The combinatorial condition is an off-by-one weak amalgamation principle, roughly stating that each walk can be refined by another walk in such a way that any two different walks refining it can be reconciled by a small perturbation. We extract the combinatorial idea at the core of the main proof of [GTZ] – that chains can be made irreconcilable by winding enough times in opposite directions – and discretize it, producing irreconcilable walks which wind around circular subgraphs.
@@
The discretization relies on a handful of classical continuum-theoretical black-boxes, all of which are stated in Section 4. In the following two sections, these black-boxes are used to derive fundamental combinatorial statements about graphs and walks on graphs. The proof of Theorem 1.2 is purely combinatorial.
@@
This approach to the study of topological and dynamical properties of compact metrizable spaces by combinatorial means traces its roots to the field of projective Fraïssé theory, first introduced in [Irwin2006]. In rough terms, one constructs an infinite compact zero-dimensional graph, the combinatorial model of $X$, and finite approximations of such graph, and studies them to derive properties about $X$, leaning on the dictionary between the combinatorics of countable structures and the dynamics of Polish non-archimedean groups put forth in the seminal paper [KPT]. The aforementioned Menger curve was studied under this lens in [PS22], and it follows from the results in [BK:Lelek], again obtained in a projective Fraïssé theoretical framework, that the Lelek fan, a one-dimensional continuum which is not locally connected, has a generic downward closed chain.
@@
Nonetheless, our strategy is innovative in that it is the first instance in which non-trivial necessary conditions for the existence of a comeager orbit for an action of a homeomorphism group of a compact connected space have been translated to a combinatorial statement. Such statement is a weakening of the weak amalgamation principle which was isolated as a criterion for the existence of comeager orbits in certain actions of automorphism groups of countable structures by Ivanov [MR1777786], and independently by Kechris and Rosendal [MR2308230]. In this paper, we restrict ourselves to finite combinatorial objects, without explicit mention of the combinatorial models of the spaces involved, though our proofs could also be used to derive the analogous statements for such combinatorial models and their groups of automorphisms.
@@
Theorem 1.2 has important dynamical consequences for the groups of homeomorphisms involved. If $G$ is a topological group, we call a continuous action $G\curvearrowright Y$ on a compact space a $G$-flow. A $G$-flow is minimal if every orbit is dense. Both classifying minimal flows of a group $G$ and, vice versa, classifying groups based on the properties of their minimal flows are central goals in topological dynamics. Some groups of homeomorphisms have very simple minimal dynamics: the only minimal flow of the group $\operatorname{Homeo}_{+}([0,1])$ of orientation preserving homeomorphisms of the interval is the trivial action and the only non-trivial minimal flow of $\operatorname{Homeo}_{+}(S^{1})$ is the action on the circle itself (see [Pestov]).
@@
A number of dividing lines regarding the complexity of the minimal dynamics of a topological group have been studied, and are particularly well understood when $G$ is Polish, i.e., separable and completely metrizable. A Polish group is extremely amenable if all of its minimal flows are trivial, it is CAP if they are metrizable (see [CAP]), and it has the generic point property if every minimal flow has a comeager orbit (see [AKL]). Every extremely amenable group is CAP, every Polish CAP group has the generic point property (see [ben2017metrizable]), and no locally compact non-compact group has the generic point property. Such properties can also be stated in terms of the universal minimal flow of $G$: CAP groups are also known as groups with metrizable universal minimal flow.
@@
Until recently (see [kaleido]), there was essentially only one known example of Polish group with the generic point property which is not CAP (see [kwiat]). In [GTZ] it is proved that the subgroup $\operatorname{Homeo}_{0}(X)$ of homeomorphisms isotopic to the identity is not CAP, for any closed surface $X$, but it was left open whether such groups have the generic point property. Since for a closed surface $X$ the flow $\operatorname{Homeo}_{0}(X)\curvearrowright\Phi(X)$ is minimal ([gutman_minimal]), Theorem 1.1 allows us to cover all but two of these cases.
@@
Let us end this introduction with a series of corollaries of Theorem 1.2. A space $X$ is homogeneous if the action $\operatorname{Homeo}(X)\curvearrowright X$ is transitive. Homogeneous Peano continua are either the circle, or a closed surface, or are locally non-planar (see [MR2674033]*Remark 8).
@@
On the other hand, the Sierpiński carpet, which is the unique one-dimensional planar Peano continuum without locally separating points (see [whyburn]), is not homogeneous. Nonetheless, it has plenty of non-locally-separating simple closed curves, since it can be constructed as the complement, in the sphere, of countably many pairwise disjoint open disks whose union is dense in the sphere itself. In particular, any open set is planar and contains a non-locally-separating simple closed curve, so the Sierpiński carpet falls within the scope of Theorem 1.2.
A one-dimensional Peano continuum with no locally separating points is either the Menger curve or has a nonempty open connected subset homeomorphic to an open subset of the Sierpiński carpet (see [MR1314981]*Lemma 1.1). Therefore
@@
We prove in Theorem 7.7 that a sufficient condition for minimality of the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$, when $X$ is a Peano continuum without locally separating points, is for $X$ to be strongly locally homogeneous, see Section 7 for the definition. For example, all closed manifolds of dimension at least $2$ are strongly locally homogeneous. Since these spaces are in particular homogeneous, they have no generic chain, by Corollary 1.5, except possibly if they are the sphere or the real projective plane. Therefore:
@@
If $X$ is a strongly locally homogeneous Peano continuum which is not the circle, the sphere, or the real projective plane, then $\operatorname{Homeo}(X)$ does not have the generic point property. In particular, its universal minimal flow is not metrizable and has no comeager orbit.
Among other examples of spaces which fit the hypotheses of Corollary 1.7 are the Menger curve, the universal $k$-dimensional compactum $\mu^{k}$, and any compact connected $\mu^{k}$-manifold, for $k\geq 1$ (see [MR0920964]*Theorem 3.2.2).
For the spaces in Corollary 1.7, we are also able to prove that the action $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$ is generically turbulent, which implies a strong non-classifiability result for the orbit equivalence relation of the action. We say that chains on $X$ are classifiable by countable structures if there is a Borel function $\mathcal{M}\colon\Phi(X)\to\mathrm{Mod}(\mathcal{L})$ to the space of structures with domain $\mathbb{N}$ in some countable relational language $\mathcal{L}$ such that
$\mathcal{C}^{\prime}\in\operatorname{Homeo}(X)\cdot\mathcal{C}\Longleftrightarrow\mathcal{M}(\mathcal{C})\cong\mathcal{M}(\mathcal{C}^{\prime}),$
@@
We remark that the first part of 1.4 could be answered in the positive by proving that chains on the sphere are classifiable by countable structures, but we believe that this is not the case.
In Appendix B, we prove that $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ is minimal, for the Sierpiński carpet $S$, and thus that $\operatorname{Homeo}(S)$ does not have the generic point property, and its universal minimal flow is not metrizable and has no comeager orbit. The proof is rather lengthy and technical; it could foreseeably be simplified by finer results on the degree of transitivity of the diagonal actions $\operatorname{Homeo}(S)\curvearrowright S^{n}$, which we were not able to find in the literature. We leave the question of classifiability by countable structures of chains on the Sierpiński carpet open.
@@ 1.1. Plan of the paper
In section 5 we construct the dictionary between open sets of chains of Peano continua and walks on graphs, and prove the required discretization results to establish the combinatorial necessary condition for the existence of a generic chain, which is Theorem 5.9. Section 6 is the heart of the paper and contains the proof of Theorem 1.2, which we obtain as a sum of Theorems 6.4, 6.6, 6.10, and 6.12, whose proofs are completely combinatorial. Section 7 is devoted to the question of minimality and generic turbulence of $\operatorname{Homeo}(X)\curvearrowright\Phi(X)$, and their consequences for classifiability of chains and topological dynamics of $\operatorname{Homeo}(X)$.
@@
Acknowledgements
We are grateful to Riccardo Camerlo, Aleksandra Kwiatkowska, Aristotelis Panagiotopoulos, and Todor Tsankov for helpful discussions at different stages of the project. We wish to thank the anonymous referees for their detailed and insightful comments.